# Horus: A Protocol for Trustless Delegation Under Uncertainty

### v5

@@ Abstract
Correctness is an emergent property of systems where exposing error is cheaper than committing it. In dynamic, low-trust environments, autonomous AI agents benefit from delegating work to sub-agents, yet correctness cannot be assured through upfront specification or centralized oversight. We propose a protocol that enforces correctness through collateralized claims in a recursive verification game. Tasks are published as intents, and solvers compete to fulfill them. Selected solvers carry out tasks under risk, with correctness checked post hoc by verifiers. Any challenger can challenge a result by staking against it to trigger the verification process. Incorrect agents are slashed and correct opposition is rewarded, with an escalation path that penalizes erroneous verifiers themselves. When incentives are aligned across solvers, challengers, and verifiers, falsification conditions make correctness the Nash equilibrium.
@@ 1 History
The twentieth century’s best response to the problem of coordination in the face of uncertainty came from Friedrich Hayek, who argued that price systems were not just resource allocators, but were distributed information processors [1]. For Hayek, price mechanisms were a way to compute solutions to coordination problems that no central planner could possibly solve. Yet, the power of price signals was constrained by the goods being sold, and as economic activity shifted away from physical goods towards services and knowledge, the limitations of markets to reliably price intangible work became increasingly clear.
@@
The advent of Bitcoin brought a practical revolution. Satoshi Nakamoto showed that it was possible to achieve consensus and settlement on facts across vast, pseudonymous networks without trusted intermediaries [7]. Vitalik Buterin extended this breakthrough by enabling not just transfers of value, but the execution of arbitrary Turing-complete logic in the form of smart contracts [8]. Digital cash, securities, and other digital assets were all enabled in this way, so long as the transaction logic was expressible and verifiable within the protocol. UMA [9] and Polymarket [10] extended these ideas by incorporating mechanisms for subjective adjudication, in which participants resolved ambiguities with economic consensus.
All of these mechanisms relied on the same underlying structure: that if exposing an error was cheaper than committing it, truth became the stable outcome. As Karl Popper observed, knowledge in complex domains depended not on exhaustive specification, but on systematic exposure to falsification [11]. UMA instantiated this principle for bounded claims, where the space of possible outcomes was discrete. However, as tasks grew more ambiguous and temporally extended, this model reached its limit. Horus builds on this work by generalizing adversarial verification to arbitrary tasks. Instead of requiring ex ante clarity or binary outcomes, it enables correctness to be adversarially surfaced, addressing the failure modes of specification and deceptive alignment identified in advanced learning systems [12].
@@ 2 Protocol
Horus enforces correctness by making error economically unprofitable. The protocol is governed by a falsification condition:
@@ 2.1 State Transitions
Figure 1: State transition diagram of the Horus protocol.
@@ 4 Deployment
Verifier pools should achieve minimum epistemic independence and diversity requirements [14]. This goes beyond purely economic decentralization: to preserve falsifiability, verifier agents must not converge on a prior that suppresses error detection. As a rule of thumb, the more bonded verifiers there are, the safer the deployment likely is.
@@ 5.4 Private Market Settlement
Private secondary markets such as EquityZen and Forge offer price discovery, yet spreads remain wide because every listing carries latent uncertainty about ownership, legal eligibility, and closing logistics [22]. Buyers cannot trust that a seller truly controls the shares, while sellers doubt that funds will arrive once paperwork clears.
@@ 6 Conclusion
We have proposed a system for delegating and verifying arbitrary tasks without relying on specification or trust. By combining collateralized execution, permissionless challenge, and recursive adjudication, Horus ensures that submitting or defending an incorrect result leads to financial loss. The protocol is robust under pseudonymity, resistant to centralized failure, and adaptive to any domain where correctness can be surfaced post hoc. Crucially, as LLM intelligence increases, the spaces where Horus can enforce correctness expand super-linearly: falsification costs $F$ trend toward zero while error detection probability $P_{e}$ approaches one, continuously strengthening the protocol’s effectiveness.

### v7

# Operator: A Protocol for Trustless Delegation Under Uncertainty

@@ Abstract
Correctness is an emergent property of systems where exposing error is cheaper than committing it. In dynamic environments, autonomous AI agents benefit from delegating work, yet correctness often cannot be assured through upfront specification or centralized oversight. We propose a protocol that enforces correctness through collateralized claims in a recursive verification game. Tasks are published as intents, and solvers compete to fulfill them. Selected solvers carry out tasks under risk, with correctness checked post hoc by verifiers. Any challenger can challenge a result by staking against it to trigger the verification process. Incorrect agents are slashed and correct opposition is rewarded, with an escalation path that penalizes erroneous verifiers themselves. When incentives are aligned across solvers, challengers, and verifiers, falsification conditions make correctness the Nash equilibrium.
@@ 1 History
The twentieth century’s best response to the problem of coordination in the face of uncertainty came from Friedrich Hayek, who argued that price systems were distributed information processors [1]. For Hayek, price mechanisms were a way to compute solutions to coordination problems that no central planner could possibly solve. Yet, the power of price signals was constrained by the goods being sold, and as economic activity shifted away from physical goods towards services and knowledge, the limitations of markets to reliably price intangible work became increasingly clear.
@@
The advent of Bitcoin brought a practical revolution. Satoshi Nakamoto showed that it was possible to achieve consensus and settlement on facts across vast, pseudonymous networks without trusted intermediaries [7]. Vitalik Buterin extended this breakthrough by enabling not just transfers of value, but the execution of arbitrary Turing-complete logic in the form of smart contracts [8]. Digital cash and other digital assets were all enabled in this way, so long as the transaction logic was expressible within the protocol. UMA [9] and Polymarket [10] extended these ideas by incorporating mechanisms that resolved ambiguities through economic consensus.
These mechanisms relied on the same underlying structure: that if exposing an error was cheaper than committing it, truth became the stable outcome. As Karl Popper observed, knowledge in complex domains depended not on exhaustive specification, but on systematic exposure to falsification [11]. UMA instantiated this principle for bounded claims, where the space of possible outcomes was discrete. However, as tasks grew more ambiguous and temporally extended, this model reached its limit. Operator builds on this work by generalizing adversarial verification to arbitrary tasks. Instead of requiring ex ante clarity or binary outcomes, it enables correctness to be adversarially surfaced, addressing the failure modes of specification and deceptive alignment identified in advanced learning systems [12].
@@ 2 Protocol
Operator enforces correctness by making error economically unprofitable. The protocol is governed by a falsification condition:
@@ 2.1 State Transitions
Figure 1: State transition diagram of the Operator protocol.
@@ 4 Deployment
Verifier agents should achieve minimum epistemic independence and diversity requirements [14]. This goes beyond purely economic decentralization: to preserve falsifiability, verifier agents must not converge on a prior that suppresses error detection. As a rule of thumb, the more bonded verifiers there are, the safer the deployment likely is.
@@ 5.4 Private Market Settlement
Private secondary markets such as EquityZen and Forge offer price discovery, but spreads remain wide because every listing carries latent uncertainty about ownership, legal eligibility, and closing logistics [22]. Buyers cannot trust that a seller truly controls the shares, while sellers doubt that funds will arrive once paperwork clears.
@@ 6 Conclusion
We have proposed a system for delegating and verifying arbitrary tasks without relying on specification or trust. By combining collateralized execution, permissionless challenge, and recursive adjudication, Operator ensures that submitting or defending an incorrect result leads to financial loss. The protocol is robust under pseudonymity, resistant to centralized failure, and adaptive to any domain where correctness can be surfaced post hoc. Crucially, as LLM intelligence increases, the spaces where Operator can enforce correctness expand super-linearly: falsification costs $F$ trend toward zero while error detection probability $P_{e}$ approaches one, continuously strengthening the protocol’s effectiveness.